# Graph Limits via Quotients

### v1

@@ Abstract
We present a new notion of limits of weighted directed graphs of growing size based on convergence of their random quotients. These limits are specified in terms of random exchangeable measures on the unit square. We call our limits grapheurs and show that these are dual to graphons in a precise sense. Grapheurs are well-suited to modeling hubs and connections between them in large graphs; previous notions of graph limits based on subgraph densities fail to adequately model such global structures as subgraphs are inherently local. Using our framework, we present an edge-based sampling approach for testing properties pertaining to hubs in large graphs. This method relies on an edge-based analog of the Szemerédi regularity lemma, whereby we show that sampling a small number of edges from a large graph approximately preserves its quotients. Finally, we observe that the random quotients of a graph are related to each other by equipartitions, and we conclude with a characterization of such random graph models.
@@ 1 Introduction
Analyzing the structure of large graphs is a challenge that arises in numerous problem domains such as bioinformatics, network tomography, operations research, and the social sciences [newman_networks, network_tomography, social_network, JACKSON2002265, bassett2017network]. Graphons and other related notions of graph limits in the literature employ subgraph densities to summarize and compare large graphs. However, subgraph-based measures are inherently local and they fail to suitably account for global structures such as hubs and connections between them, which are commonly encountered in real-world networks [OKELLY1998171, hub_and_spoke, power2013evidence, hubs_in_drugs, van2013network]. For example, most subgraphs of a star, the prototypical example of a hub, do not contain any edges.
@@
Here we view $\rho(f_{k,n}):\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{k\times k}$ as a ‘linear representation’ of $f_{k,n}$. In words, the vertices of the quotient $\rho(f_{k,n})G$ correspond to the parts in the partition defined by $f_{k,n}$, and the edge weight between parts $i,j\in[k]$ is given by sums of all edge weights in $G$ between vertices in parts $f_{k,n}^{-1}(i)$ and $f_{k,n}^{-1}(j)$. Thus, a quotient of $G$ represents the edge weights between each pair of parts in a vertex partition.
@@
As an illustration, we give in Figures 1-2 examples of small quotients of two networks. The first, depicted in Figure 1, is a traffic network representing the fraction of trips starting and ending at each zone of the Gold Coast, Australia [traffic_networks]. The second, depicted in Figure 2, is a brain connectome, the network of connections between neurons of an adult male C. elegans roundworm [cook2019whole, worms_brain]. The structure of each quotient indicates some of the features of the underlying large network. As we will see in Section 4.3, the presence of hubs in the original network leads to nonuniform distribution of edge weight in their random quotients.
@@
Figure 1: Traffic network for the Gold Coast, Australia, and an associated random quotient. The city is divided into 1068 zones, and the weight of an edge is the fraction of trips from a dataset with given start and end zones [traffic_networks]. Here we plot the first 300 of these zones and their interconnections, displaying the width of an edge proportionally to its weight. We color the vertices according to their image in the quotient, and color and edge connecting two vertices with the same image by the color of that image. The nonuniformity in the edge weights of the quotient indicates the presence of hubs that we can see in the plot of the entire network.
@@
Figure 2: Network of chemical connections among the 575 neurons of adult male C. elegans (a species of roundworm), and its random quotient. Edges are directed from pre-synaptic to post-synaptic cells, and their weight is the total number of EM serial sections of connectivity [cook2019whole, worms_brain]. The nonuniform distribution of self-loops and the unequal in-degrees and out-degrees among the four vertices in the quotient again indicate the presence of hubs.
@@
In this paper, we define a new notion of convergence for sequences of growing-sized graphs based on convergence of their random quotients. In analogy with graphons [lovasz2012large, §1], we define suitable limit objects in terms of random measures for any quotient-convergent sequence of graphs. Finally, we show that finite (random) graphs sampled from our limit objects are quotient-convergent to the limit almost surely. The key concepts in our paper are dual to those in the literature on graphons, and we emphasize this duality throughout our development. In particular, we shall see that random quotients and random subgraphs, in terms of which graphon convergence is defined, are random linear maps which are adjoints of each other. Further, a dual analog of the cut metric plays a prominent role in our development, and we derive a dual Szemerédi lemma as a consequence of our results. We describe next these contributions in more detail.
@@ 1.1 Our Contributions
To set the stage, we define a random quotient of a graph on $n$ vertices in terms of a uniformly random map $F_{k,n}\colon[n]\to[k]$. That is, each $i\in[n]$ gets sent independently to a uniform element $F_{k,n}(i)\in[k]$. Equivalently, this is a uniformly random ordered partition of $[n]$ into $k$ parts, given by the sequence of (random) fibers $F_{k,n}^{-1}(j)$ for $j\in[k]$. We denote by $\Delta^{n\times n}$ the set of $n\times n$ matrices with nonnegative entries that sum to one.
@@
Note that $\rho(F_{n,k})^{\star}G_{n}$ is the induced subgraph on $k$ vertices sampled with replacement from the vertex set $[n]$ of $G_{n}$, in terms of which graphon convergence is defined [lovasz2012large, §9], giving a first sense in which our theory is dual to that of graphons. We remark that there is an alternative view of graphon limits via convergence of quotients [lovasz2012large, §12], but these latter quotients differ from ours. In particular, these latter quotients cannot detect hubs in a large graph, as we explain in Section 1.2 below.
@@ 1.1.1 Combinatorial Aspects of Quotient Convergence (Section 2)
As our first contribution, we analyze combinatorial aspects of quotient convergence in Section 2. Whereas previous notions of graph convergence were based on convergence of local “motifs” in the form of homomorphism densities (appropriately normalized), we show that our quotient convergence is equivalent to convergence of homomorphism numbers. More precisely, for a multigraph $H\in\mathbb{N}^{k\times k}$ possibly containing parallel edges and self-loops, and a general weighted graph $G\in\mathbb{R}^{n\times n}$ on $n$ vertices, the homomorphism number and homomorphism density of $H$ in $G$ are given by [lovasz2012large, Eq. (5.3)]
@@
The terminology comes from the fact that when $H$ and $G$ are simple graphs, then $\mathrm{hom}(H;G)$ is the number of graph homomorphisms from $H$ to $G$ and $t(H;G)$ counts the average number of times $H$ occurs in a uniformly random subgraph of $G$. In particular, convergence of the homomorphism densities $t(H;\cdot)$ is used to define limits of dense graphs [LOVASZ2006933, convergent_seqs1].
@@
The role of homomorphism densities in graphons is played by quotient densities in our framework. We prove that a sequence of graphs is quotient-convergent if all of their quotient densities converge.
@@
The proof of this result is given in Section 2 based on expressing quotient densities in terms of (injective) homomorphism numbers. This expression involves a partial order on the collection of multigraphs induced by the quotient operation.
@@ 1.1.2 Limit Objects for Quotient Convergence (Section 3)
Our second contribution consists of constructing analytic limit objects together with a notion of distance between them that allows us to compare graphs of different sizes, akin to graphons and the cut metric for dense graph sequences. Note that appending isolated vertices to a graph $G\in\Delta^{n\times n}$ or relabeling its vertices does not change the distribution of its quotients $\rho(F_{k,n})G_{n}$ for each $k\in\mathbb{N}$. Therefore, the limit object we associate to $G$ must be the same as that associated to any other graph differing from it by extra isolated vertices and vertex relabeling. We describe our limit objects in terms of certain random exchangeable measures on the square, which we define next.
@@
Note that relabeling the vertices of $G$ or appending isolated vertices to it yields the same random measure in this way. More generally, Kallenberg showed that all random exchangeable measures on $0,1]^{2}$ have the form [kallenberg1990exchangeable]:
@@
where $(i)$ $\lambda$ is the uniform measure on $[0,1]$, $\lambda^{2}$ is the uniform measure on $0,1]^{2}$, and $\lambda_{D}$ is the uniform measure on the diagonal $\{(x,x):x\in[0,1]\}$; $(ii)$ $T_{1},T_{2},\ldots\overset{iid}{\sim}\mathrm{Unif}([0,1])$; and $(iii)$ $E_{i,j},\sigma_{i},\varsigma_{i},\theta,\vartheta\geq 0$ may potentially be random but must be independent of the $(T_{i})$ and must satisfy $\sum_{i,j}E_{i,j}+\sum_{i}(\sigma_{i}+\varsigma_{i})+\theta+\vartheta=1$ almost surely. Moreover, the collection of extremal random exchangeable measures, i.e., those that are not mixtures of other such random measures, are precisely the ones for which the coefficients $E,\sigma,\varsigma,\theta,\vartheta$ in (5) can be taken to be deterministic.
@@
The coefficients $(E,\sigma,\varsigma,\theta,\vartheta)$ describing a fixed grapheur $\mathsf{M}\in\mathcal{M}$ are essentially unique. Specifically, note that for any permutation $\pi\colon\mathbb{N}\to\mathbb{N}$, the coefficients $(E,\sigma,\varsigma,\theta,\vartheta)$ and $(\pi E\pi^{\top},\pi\sigma,\pi\varsigma,\theta,\vartheta)$ define the same random exchangeable distribution via (5). Further, adding zero rows to $E$ and corresponding zero entries to $\sigma$, or zero columns to $E$ and corresponding zero entries to $\varsigma$, does not change the resulting random measure. Kallenberg showed in [kallenberg1990exchangeable, Prop. 3] that these are the only possible ambiguities in the coefficients. In particular, note that $\mathsf{M}_{G}\in\mathcal{M}$ is a grapheur for any finite graph $G$, and that $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ for two such graphs $G,G^{\prime}$ if and only if they differ by isolated vertices and vertex relabeling.
@@
    Any sequence of graphs $(G_{n}\in\Delta^{n\times n})$ has a quotient-convergent subsequence.
@@
$t_{Q}(H;\mathsf{M})=\mathbb{E}\prod_{i,j=1}^{k}(\mathsf{G}_{k}[\mathsf{M}])_{i,j}^{H_{i,j}}=\int\prod_{i,j\in[k]}\mu(I_{i}^{(k)}\times I_{j}^{(k)})^{H_{i,j}}\,d\mathsf{M}(\mu),$ (9)
for any multigraph $H\in\mathbb{N}^{k\times k}$, analogously to the extension of homomorphism densities to graphons [lovasz2012large, §7.2]. In terms of quotient densities, the sequence $(G_{n})$ converges to the grapheur $\mathsf{M}$ if and only if $t_{Q}(H;G_{n})\to t_{Q}(H;\mathsf{M})$ for all multigraphs $H$, see Section 3.1.
Having characterized our limit objects, we interpret in Section 3.2 each of the parameters $E,\sigma,\varsigma,\theta,\vartheta$ describing a grapheur (5) in terms of certain features of the large graphs converging to it. Informally, the array $E\in\mathbb{R}_{\geq 0}^{\mathbb{N}\times\mathbb{N}}$ specifies the connections between a discrete collection of hubs, with each entry corresponding to an edge with an asymptotically positive fraction of the total edge weight. The arrays $\sigma,\varsigma\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ specify directed stars centered at hubs, with each entry corresponding to the fraction of out-degree or in-degree of a hub from all the edges incident on it with asymptotically vanishing edge weights. Lastly, the parameters $\theta$ and $\vartheta$ specify the remaining edge weight that comes from all edges between vertices with asymptotically vanishing degrees and their self-loops. In particular, we say that a grapheur $\mathsf{M}$ does not have hubs if $E=0$ and $\sigma=\varsigma=0$, and we say that a sequence $(G_{n})$ of graphs does not contain hubs asymptotically if it quotient-converges to such a grapheur.
Finally, we expand on the duality between graphons and grapheurs. We have seen previously that the sampling operators used to define each notion of graph limit, namely, those extracting random subgraphs and random quotients, respectively, are duals of each other. We further show in Section 3.3 that the space of graphons endowed with the cut metric is dual to the space $\mathcal{M}$ endowed with the $W_{\square}$ metric (7) in a precise sense.
@@ 1.1.3 Edge Sampling and Property Testing (Section 4)
As our next contribution, we apply our framework to test properties of large graphs based on small summaries of them. Specifically, we focus on testing properties given by graph parameters that are continuous with respect to quotient convergence. While it would be natural to test such properties using random quotients, forming these requires access to the entire large graph, which limits their utility in practice. Motivated by this challenge, we develop an edge-based sampling perspective on quotient-convergence in Section 4.1. Specifically, we sample edges from a graph proportionally to their edge weights to form random graphs $\mathsf{G}^{(n)}$ with $n$ edges, and extend this sampling procedure to any grapheur $\mathsf{M}\in\mathcal{M}$. We detail the general sampling procedure in Section 4.1.
We prove the following convergence results for the sequence of random graphs $(\mathsf{G}^{(n)})$ sampled from a grapheur $\mathsf{M}$ back to $\mathsf{M}$, which are edge-based analogs of the second sampling lemma and Szemerédi’s regularity lemma for graphons [lovasz2012large, Lemma 10.16, Lemma 9.15]. Here and throughout the paper, a superscript $G^{(n)}$ denotes a graph on $n$ edges (and hence at most $2n$ vertices) while a subscript $G_{n}$ denotes a graph on $n$ vertices.

### v2

@@ Abstract
We present a new notion of limits of weighted directed graphs of growing size based on convergence of their random quotients. These limits are specified in terms of random exchangeable measures on the unit square. We call our limits grapheurs and show that these are dual to graphons in a precise sense. Grapheurs are well-suited to modeling global structure in large graphs such as hubs and connections between them; previous notions of graph limits based on subgraph densities fail to adequately model such global structure as subgraphs are inherently local. Using our framework, we characterize properties of large graphs that are continuous with respect to our limits and present an edge-based sampling approach for testing them. This method relies on an edge-based analog of the Szemerédi regularity lemma, whereby we show that sampling a constant number of edges from an arbitrarily-large graph approximately preserves its quotients. Finally, we observe that the random quotients of a graph are related to each other by equipartitions, and we conclude with a characterization of such random graph models.
@@ 1 Introduction
Analyzing the structure of large graphs is a challenge that arises in numerous problem domains such as bioinformatics, network tomography, operations research, and the social sciences [36, 13, 42, 22, 2]. Graphons and other related notions of graph limits in the literature employ subgraph densities to summarize and compare large graphs. However, subgraph-based measures are inherently local and they fail to suitably account for global structures such as hubs and connections between them, which are commonly encountered in real-world networks [37, 9, 40, 21, 43]. For example, most subgraphs of a star, the prototypical example of a hub, do not contain any edges.
@@
Here we view $\rho(f_{k,n}):\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{k\times k}$ as a ‘linear representation’ of $f_{k,n}$. In words, the vertices of the quotient $\rho(f_{k,n})G$ correspond to the parts in the partition defined by $f_{k,n}$, and the edge weight between parts $i,j\in[k]$ is given by the sum of all edge weights in $G$ between vertices in $f_{k,n}^{-1}(i)$ and $f_{k,n}^{-1}(j)$. Thus, a quotient of $G$ represents the edge weights between each pair of parts in a vertex partition.
@@
As an illustration, we give in Figures 1-2 examples of small quotients of two networks. The first, depicted in Figure 1, is a traffic network representing the fraction of trips starting and ending at each zone of the Gold Coast, Australia [41]. The second, depicted in Figure 2, is a brain connectome, the network of connections between neurons of an adult male C. elegans roundworm [14, 35]. The structure of each quotient indicates some of the features of the underlying large network. As we will see in Section 4.3, the presence of hubs in the original network leads to nonuniform distribution of edge weight in their random quotients.
@@
Figure 1: Traffic network for the Gold Coast, Australia, and an associated random quotient. The city is divided into 1068 zones, and the weight of an edge is the fraction of trips from a dataset with given start and end zones [41]. Here we plot the first 300 of these zones and their interconnections, displaying the width of an edge proportionally to its weight. We color the vertices according to their image in the quotient, and color an edge connecting two vertices with the same image by the color of that image. The nonuniformity in the edge weights of the quotient indicates the presence of hubs that we can see in the plot of the entire network.
@@
Figure 2: Network of chemical connections among the 575 neurons of adult male C. elegans (a species of roundworm), and a random quotient. Edges are directed from pre-synaptic to post-synaptic cells, and their weight is the total number of EM serial sections of connectivity [14, 35]. The nonuniform distribution of self-loops and the unequal in-degrees and out-degrees among the four vertices in the quotient again indicate the presence of hubs.
@@
In this paper, we define a new notion of convergence for sequences of growing-sized graphs based on convergence of their random quotients. In analogy with graphons [34, §1], we define suitable limit objects in terms of random measures for any quotient-convergent sequence of graphs. Finally, we show that finite (random) graphs sampled from our limit objects are quotient-convergent to the limit almost surely. The key concepts in our paper are dual to those in the literature on graphons, and we emphasize this duality throughout our development. In particular, we shall see that random quotients and random subgraphs, in terms of which graphon convergence is defined, are random linear maps which are adjoints of each other. Further, a dual analog of the cut metric plays a prominent role in our development, and we derive an edge-based analog of the Szemerédi regularity lemma as a consequence of our results. We describe next these contributions in more detail.
@@ 1.1 Our Contributions
To set the stage, we define a random quotient of a graph on $n$ vertices in terms of a uniformly random map $F_{k,n}\colon[n]\to[k]$ sending each $i\in[n]$ independently to a uniformly-random element $F_{k,n}(i)\in[k]$. Equivalently, this is a uniformly random ordered partition of $[n]$ into $k$ parts, given by the sequence of (random) fibers $F_{k,n}^{-1}(j)$ for $j\in[k]$. We denote by $\Delta^{n\times n}$ the set of $n\times n$ matrices with nonnegative entries that sum to one.
@@
Denoting by $\rho(F_{n,k})^{\star}\colon\mathbb{R}^{n\times n}\to\mathbb{R}^{k\times k}$ the adjoint of $\rho(F_{n,k})$ with respect to the Frobenius inner product, note that $\rho(F_{n,k})^{\star}G_{n}$ is the induced subgraph on $k$ vertices sampled with replacement from the vertex set $[n]$ of $G_{n}$, in terms of which graphon convergence is defined [34, §9]. This is a first sense in which our theory is dual to that of graphons. We remark that there is an alternative view of graphon limits via convergence of quotients [34, §12], but these latter quotients differ from ours. In particular, these latter quotients cannot detect hubs in a large graph, as we explain in Section 1.2 below.
@@ 1.1.1 Combinatorial Aspects of Quotient Convergence (Section 2)
As our first contribution, we analyze combinatorial aspects of quotient convergence in Section 2. Whereas previous notions of graph convergence were based on convergence of local “motifs” in the form of homomorphism densities (appropriately normalized), we show that our quotient convergence is equivalent to convergence of homomorphism numbers. More precisely, for a multigraph $H\in\mathbb{N}^{k\times k}$ possibly containing parallel edges and self-loops, and a general weighted graph $G\in\mathbb{R}^{n\times n}$ on $n$ vertices, the homomorphism number and homomorphism density of $H$ in $G$ are given by [34, Eq. (5.3)]
@@
The terminology comes from the fact that when $H$ and $G$ are simple graphs, then $\mathrm{hom}(H;G)$ is the number of graph homomorphisms from $H$ to $G$ and $t(H;G)$ counts the average number of times $H$ occurs in a uniformly random subgraph of $G$. In particular, convergence of the homomorphism densities $t(H;\cdot)$ is used to define limits of dense graphs [32, 4].
@@
The role of homomorphism densities in graphon theory is played by quotient densities in our framework. We prove that a sequence of graphs is quotient-convergent if all of their quotient densities converge.
@@
The proof of this result is given in Section 2, and is based on expressing quotient densities in terms of (injective) homomorphism numbers. This expression involves a partial order on the collection of multigraphs induced by the quotient operation.
@@ 1.1.2 Limit Objects for Quotient Convergence (Section 3)
Our second contribution consists of constructing analytic limit objects for quotient-convergent sequences together with a notion of distance between them that allows us to compare graphs of different sizes, akin to graphons and the cut metric for dense graph sequences. Note that appending isolated vertices to a graph $G\in\Delta^{n\times n}$ or relabeling its vertices does not change the distribution of its quotients $\rho(F_{k,n})G_{n}$ for each $k\in\mathbb{N}$. Therefore, the limit object we associate to $G$ must be the same as that associated to any other graph differing from it by extra isolated vertices and vertex relabeling. We describe our limit objects in terms of certain random exchangeable measures on the square, which we define next.
@@
Note that relabeling the vertices of $G$ or appending isolated vertices to it yields the same random measure in this way. More generally, Kallenberg showed that all random exchangeable measures on $0,1]^{2}$ have the form [28]:
@@
where $(i)$ $\lambda$ is the uniform measure on $[0,1]$, $\lambda^{2}$ is the uniform measure on $0,1]^{2}$, and $\lambda_{D}$ is the uniform measure on the diagonal $\{(x,x):x\in[0,1]\}$; $(ii)$ $T_{1},T_{2},\ldots\overset{iid}{\sim}\mathrm{Unif}([0,1])$; and $(iii)$ $E_{i,j},\sigma_{i},\varsigma_{i},\theta,\vartheta\geq 0$ may potentially be random but must be independent of the $(T_{i})$ and must satisfy $\sum_{i,j}E_{i,j}+\sum_{i}(\sigma_{i}+\varsigma_{i})+\theta+\vartheta=1$ almost surely. Moreover, the collection of extremal random exchangeable measures, i.e., those that are not mixtures of other such random measures, are precisely the ones for which the coefficients $E,\sigma,\varsigma,\theta,\vartheta$ in (5) can be taken to be deterministic. We call such extremal random measures grapheurs, and show that these are precisely limits of quotient-convergent graph sequences.
@@
The coefficients $(E,\sigma,\varsigma,\theta,\vartheta)$ describing a fixed grapheur $\mathsf{M}\in\mathcal{M}$ are essentially unique. Specifically, note that for any permutation $\pi\colon\mathbb{N}\to\mathbb{N}$, the coefficients $(E,\sigma,\varsigma,\theta,\vartheta)$ and $(\pi E\pi^{\top},\pi\sigma,\pi\varsigma,\theta,\vartheta)$ define the same random exchangeable distribution via (5). Further, adding zero rows to $E$ and corresponding zero entries to $\sigma$, or zero columns to $E$ and corresponding zero entries to $\varsigma$, does not change the resulting random measure. Kallenberg showed in [28, Prop. 3] that these are the only possible ambiguities in the coefficients. In particular, note that $\mathsf{M}_{G}\in\mathcal{M}$ is a grapheur for any finite graph $G$, and that $\mathsf{M}_{G}=\mathsf{M}_{G^{\prime}}$ for two such graphs $G,G^{\prime}$ if and only if they differ by isolated vertices and vertex relabeling.
@@
    The space of grapheurs $\mathcal{M}$ is compact. In particular, any sequence of graphs $(G_{n}\in\Delta^{n\times n})$ has a quotient-convergent subsequence.
@@
$t_{Q}(H;\mathsf{M})=\mathbb{E}\prod_{i,j=1}^{k}(\mathsf{G}_{k}[\mathsf{M}])_{i,j}^{H_{i,j}}=\int\prod_{i,j=1}^{k}\mu(I_{i}^{(k)}\times I_{j}^{(k)})^{H_{i,j}}\,d\mathsf{M}(\mu),$ (9)
for any multigraph $H\in\mathbb{N}^{k\times k}$, analogously to the extension of homomorphism densities to graphons [34, §7.2]. In terms of quotient densities, the sequence $(G_{n})$ converges to the grapheur $\mathsf{M}$ if and only if $t_{Q}(H;G_{n})\to t_{Q}(H;\mathsf{M})$ for all multigraphs $H$, see Section 3.1.
Having characterized our limit objects, we interpret in Section 3.2 each of the parameters $E,\sigma,\varsigma,\theta,\vartheta$ describing a grapheur (5) in terms of certain features of the large graphs converging to it. Informally, the array $E\in\mathbb{R}_{\geq 0}^{\mathbb{N}\times\mathbb{N}}$ specifies the connections between a discrete collection of hubs, with each entry corresponding to an edge with an asymptotically positive fraction of the total edge weight. The arrays $\sigma,\varsigma\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ specify directed stars centered at hubs, with each entry corresponding to the fraction of out-degree or in-degree of a hub from all the edges incident to it with asymptotically vanishing edge weights. Lastly, the parameters $\theta$ and $\vartheta$ specify the remaining edge weight that comes from all edges between vertices with asymptotically vanishing degrees and their self-loops. In particular, we say that a grapheur $\mathsf{M}$ does not have hubs if $E=0$ and $\sigma=\varsigma=0$, and we say that a sequence $(G_{n})$ of graphs does not contain hubs asymptotically if it quotient-converges to such a grapheur.
Finally, we expand on the duality between graphons and grapheurs. We have seen previously that the sampling operators used to define each notion of graph limit, namely, those extracting random subgraphs and random quotients, respectively, are duals of each other. We further show in Section 3.3 that the space of graphons endowed with the cut metric is dual to the space $\mathcal{M}$ endowed with the $W_{\square}$ metric (7) as follows.
@@ 1.1.3 Edge Sampling and Property Testing (Section 4)
As our next contribution, we apply our framework to test properties of large graphs based on small summaries of them. Specifically, we focus on testing properties given by graph parameters that are continuous with respect to quotient convergence, including quotient densities, homomorphism numbers, clustering and centrality coefficients, see Section 4.3. While it would be natural to test such properties using random quotients, forming these requires access to the entire large graph, which limits their utility in practice. Motivated by this challenge, we develop an edge-based sampling perspective on quotient-convergence in Section 4.1. Specifically, we sample edges from a graph proportionally to their edge weights to form random graphs $\mathsf{G}^{(n)}$ with $n$ edges, and extend this sampling procedure to any grapheur $\mathsf{M}\in\mathcal{M}$. We detail the general sampling procedure in Section 4.1.
We prove the following convergence results for the sequence of random graphs $(\mathsf{G}^{(n)})$ sampled from a grapheur $\mathsf{M}$ back to $\mathsf{M}$, which are edge-based analogs of the second sampling lemma and Szemerédi’s regularity lemma for graphons [34, Lemma 10.16, Lemma 9.15]. Here and throughout the paper, a superscript $G^{(n)}$ denotes a graph on $n$ edges (and hence at most $2n$ vertices) while a subscript $G_{n}$ denotes a graph on $n$ vertices.